OpenStudy (anonymous):

How to find the integral of 3y^3e^3xy-xdx

geerky42 (geerky42):

Like this? \(\Large \displaystyle \int\limits \left(3y^3e^3xy-x \right)\text{d}x\)

OpenStudy (anonymous):

E to the 3xy

OpenStudy (anonymous):

Raised to

OpenStudy (anonymous):

And minus 1 not x sorry

OpenStudy (anonymous):

@geerky42 Any idea?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+3y%5E3e%5E3xy-xdx Hopefully that can help you out

OpenStudy (anonymous):

Yeah but I need to know how to get there

OpenStudy (anonymous):

actually I think this is what you wanted http://www.wolframalpha.com/input/?i=integral+of+%283y%5E3*e%5E%283xy%29-x%29+dx

OpenStudy (anonymous):

The integral goes from y^3 to y^2 somehow

OpenStudy (anonymous):

Thats the part I don't get

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/DE/Exact.aspx example 5

OpenStudy (anonymous):

Hum Unfortunately im not sure on that one sorry

OpenStudy (anonymous):

Thats the answwer

OpenStudy (anonymous):

I jus don't understand how they got it